# In-situ tunable interaction with an invertible sign between a fluxonium and a post cavity

### v1

@@ Abstract
Quantum computation with bosonic modes presents a powerful paradigm for harnessing the principles of quantum mechanics to perform complex information processing tasks. In constructing a bosonic qubit with superconducting circuits, nonlinearity is typically introduced to a cavity mode through an ancillary two-level qubit. However, the ancilla’s spurious heating has impeded progress towards fully fault-tolerant bosonic qubits. The ability to in-situ decouple the ancilla when not in use would be beneficial but has not been realized yet. This work presents a novel architecture for quantum information processing, comprising a 3D post cavity coupled to a fluxonium ancilla via a readout resonator. This system’s intricate energy level structure results in a complex landscape of interactions whose sign can be tuned in situ by the magnetic field threading the fluxonium loop. Our results could significantly advance the lifetime and controllability of bosonic qubits.
@@
Superconducting microwave cavities are a great candidate to host bosonic qubits as their lifetimes are enhanced by at least an order of magnitude compared to conventional junction-based superconducting qubits, such as transmons and fluxonia. To control such a bosonic mode, a nonlinearity is typically introduced to the cavity through a transmon qubitKoch et al. (2007), referred to as ancilla. However, it has been shown that this ancilla is the main limitation of the bosonic qubit due to residual excitations out of the computational space and its chaotic behavior under strong drivesCohen et al. (2023). This motivated a search for alternatives: removing the transmon overallRéglade et al. (2024) or using an ancilla bosonic qubit Royer, Singh, and Girvin (2020). While both approaches succeed in their objective, they come with disadvantages. The former uses an increased control complexity, i.e. a bigger instrument control stack is needed, while the latter suffers from weaker interaction strengths leading to longer experiment times. A perfect replacement would be an ancilla that can be decoupled selectively from the bosonic mode after the state preparation and tomography.
@@ III Experimental Results
The extracted data is plotted together with the numerical model predictions in Fig. 3. We notice that in Fig. 3b, the dispersive shift $\chi_{QC}$ is smoothly tunable between positive and negative values through zero, meaning we can switch off the interaction between the fluxonium and the cavity. The largest negative dispersive shift we measured is $\text{max}(\chi_{QC})/2\pi=-51\pm 2$kHz, which is comparable to other recent works Eickbusch et al. (2022); Sivak et al. (2020); Ni et al. (2023). Also, the closest to zero we measured is $\text{min}(\chi_{QC})/2\pi=900\pm 350$ Hz. While such measured results can also be achieved with a tunable transmon as an ancillaValadares et al. (2024), in this setup it is possible to tune the interaction strictly to zero. Furthermore, we can tune to the opposite sign of $\chi_{QC}$, with our largest measured value at $78\pm 9$ kHz, similar to tunable transmon implementations using a complex inductive couplerChen et al. (2014).
@@ IV Conclusion and Outlook
To summarize, we have demonstrated that a fluxonium is a viable ancilla to a high-Q coaxial cavity. We explore for the first time this tunable interaction with an invertible sign, which enables in-situ control strategies for building bosonic codes. The fact that the ancilla can be decoupled from the bosonic mode alleviates the stringent requirement for a high-coherence ancilla without requiring a trade-off with the bosonic qubit control capabilities.
@@ Appendix C Simulations
Firstly, all interaction terms are required as rotating wave approximation is not applicable in order to get the correct interaction strengths. Secondly, the higher cavity modes cannot be neglected and will influence the bare frequencies of the harmonic oscillators, needed for the fit. They will also push the qubit resonance to lower frequencies than anticipated. In Fig. 6 we show the influence of higher cavity modes on the qubit fundamental frequency. Three higher modes of the cavity are included with direct coupling to the resonator, the highest frequency mode also is coupled directly to the qubit. The coupling strengths were based on expectation values from the scaling with $\sqrt{\omega}$ and then further increased to amplify their effect. As the frequencies of those modes were never measured, but extracted from simulations, the actual cavity modes could coincide with the qubit levels, which would cause a further shift in the transition $f_{01}$. Such a coincidence could explain why around $\Phi_{\text{ext}}/\Phi_{0}=0$ the measured qubit frequencies fit better to the cavity mode at $f_{c}^{1}$.

### v2

@@ Abstract
Quantum computation with bosonic modes presents a powerful paradigm for harnessing the principles of quantum mechanics to perform complex information processing tasks. In constructing a bosonic qubit with superconducting circuits, nonlinearity is typically introduced to a cavity mode through an ancillary two-level qubit. However, the ancilla’s spurious heating has impeded progress towards fully fault-tolerant bosonic qubits. The ability to in situ decouple the ancilla when not in use would be beneficial but has so far only been realized with tunable couplers or additional parametric drives. This work presents a novel architecture for quantum information processing, comprising a 3D post cavity coupled to a fluxonium ancilla via a readout resonator. This system’s intricate energy level structure results in a complex landscape of interactions whose sign can be tuned in situ by the magnetic field threading the fluxonium loop without the need of additional elements. Our results could significantly advance the lifetime and controllability of bosonic qubits.
@@
Superconducting microwave cavities are a great candidate to host bosonic qubits as their lifetimes are enhanced by at least an order of magnitude compared to conventional junction-based superconducting qubits, such as transmons and fluxonia. To control such a bosonic mode, a nonlinearity is typically introduced to the cavity through a transmon qubitKoch et al. (2007), referred to as ancilla. However, it has been shown that this ancilla is the main limitation of the bosonic qubit due to residual excitations out of the computational space and its chaotic behavior under strong drivesCohen et al. (2023). This motivated a search for alternatives: removing the transmon overallRéglade et al. (2024) or using an ancilla bosonic qubit Royer, Singh, and Girvin (2020). While both approaches succeed in their objective, they come with disadvantages. The former uses an increased control complexity, i.e. a bigger instrument control stack is needed, while the latter suffers from weaker interaction strengths leading to longer experiment times. A perfect replacement would be an ancilla that can be decoupled selectively from the bosonic mode after the state preparation and tomography. So far, this has been realized by introducing either additional parametric drivesRosenblum et al. (2018) or elements such as tunable couplersChen et al. (2014) or a combination of bothNoh et al. (2023).
@@ III Experimental Results
The extracted data is plotted together with the numerical model predictions in Fig. 3. We notice that in Fig. 3b, the dispersive shift $\chi_{QC}$ is smoothly tunable between positive and negative values through zero, meaning we can switch off the interaction between the fluxonium and the cavity. The largest negative dispersive shift we measured is $\text{max}(\chi_{QC})/2\pi=-51\pm 2$kHz, which is comparable to other recent works Eickbusch et al. (2022); Sivak et al. (2023); Ni et al. (2023). Also, the closest to zero we measured is $\text{min}(\chi_{QC})/2\pi=900\pm 350$ Hz. While such measured results can also be achieved with a tunable transmon as an ancillaValadares et al. (2024), in this setup it is possible to tune the interaction strictly to zero. Furthermore, we can tune to the opposite sign of $\chi_{QC}$, with our largest measured value at $78\pm 9$ kHz, similar to tunable transmon implementations using additional elements like inductive couplerChen et al. (2014); Noh et al. (2023).
@@ IV Conclusion and Outlook
To summarize, we have demonstrated that a fluxonium is a viable ancilla to a high-Q coaxial cavity. We explore for the first time that this platform realizes tunable interaction with an invertible sign without using supplementary circuit elements like tunable couplers or applying parametric drives. This appends another element to the toolbox to control higher coherence microwave cavities, which could enable novel control strategies for building bosonic codes. The fact that the ancilla can be decoupled from the bosonic mode alleviates the stringent requirement for a high-coherence ancilla without requiring a trade-off with the bosonic qubit control capabilities.
@@ Appendix C Simulations
Firstly, all interaction terms are required as rotating wave approximation is not applicable in order to get the correct interaction strengths. Secondly, the higher cavity modes cannot be neglected and will influence the bare frequencies of the harmonic oscillators, needed for the fit. They will also push the qubit resonance to lower frequencies than anticipated. In Fig. 6 we show the influence of higher cavity modes on the qubit fundamental frequency. Three higher modes of the cavity are included with direct coupling to the resonator, the highest frequency mode also is coupled directly to the qubit. The coupling strengths were based on expectation values from the scaling with $\sqrt{\omega}$ and then further increased to amplify their effect. As the frequencies of those modes were never measured, but extracted from simulations, the actual cavity modes could coincide with the qubit levels, which would cause a further shift in the transition $f_{01}$. Such a coincidence could explain why around $\Phi_{\text{ext}}/\Phi_{0}=0$, the measured qubit frequencies are closer to the cavity mode $f_{c}^{1}$ and not to the simulated frequencies for $f_{01}$.